# radarODE-MTL: A Multi-Task Learning Framework with Eccentric Gradient Alignment for Robust Radar-Based ECG Reconstruction

### v1

@@ I Introduction
  • To the best of our knowledge, this is the first work that models the radar-based long-term ECG recovery as three tasks and realized by an end-to-end MTL framework named radarODE-MTL, improving the robustness against both constant and abrupt noises.
  • Sufficient experiments show that the proposed radarODE-MTL with EGA optimization strategy outperforms other frameworks and optimization strategies under various noise conditions and datasets.
@@ II-A1 Signal Model for Cardiac Mechanical Activities
The first step for any contactless vital sign monitoring is to obtain the displacement near the chest region by unwrapping the chest displacement $\tilde{x}(t)$ from the phase change $\Delta\phi(t)$ of the received radar signal with wavelength $\lambda$ as:
$\Delta\phi(t)=\frac{4\pi\tilde{x}(t)}{\lambda}$ (1)
The demodulated chest displacement contains the interested cardiac mechanical activities with multiple noises, such as respiration noise [17] and RBM noise [2]. In ECG recovery, some periodic or high-frequency noise can be filtered in pre-processing following the methods in [6], while some constant or abrupt noises are hard to remove. Therefore, the original chest displacement $\tilde{x}(t)$ can be processed to $x(t)$ for ECG recovery after initial noise removal as:
$x(t)=x_{c}(t)+n_{abr}(t)+n_{con}(t)$ (2)
where $x_{c}(t)$ contains the interested cardiac mechanical activities, $n_{abr}(t)$ represents the abrupt noises (e.g., RBM) and $n_{con}(t)$ describes many other constant noises that affect the signal-to-noise ratio (SNR), such as thermal noise [6, 18], monitoring from random directions [19] and long-range monitoring [17].
According to our previous work [5], the fine-grained cardiac mechanical activities include aortic valve opening/closure (AO/AC) and mitral valve opening/closure (MO/MC), revealed by the corresponding prominent vibrations $v_{1}$ and $v_{2}$ as measured in radar signal $x(t)$ as depicted in Figure 1. Therefore, the signal model $x(t)$ can be further refined into subtle cardiac activities for $K$ cardiac cycles as:
@@
where $a_{1}^{k}$, $b_{1}^{k}$ and $a_{2}^{k}$, $b_{2}^{k}$ jointly determine the amplitudes and lengths of the first and second prominent vibrations for $k^{th}$ cardiac cycle, $f_{1}^{k}$, $f_{2}^{k}$ are the corresponding central frequencies and $T_{1}^{k}$, $T_{2}^{k}$ represent when the vibrations happen.
@@ II-A2 Model of Domain Transformation
  • Task $1$: The reconstruction of the morphological features aims to learn the mapping function $\mathcal{T}(\cdot)$ only for the single-cycle cardiac activities from $x(t)$ to the ECG pieces as $x_{ecg}(t)=\mathcal{T}(x(t))$.
@@
  • Task $3$: The prediction of the length of a single cardiac cycle is equivalent to finding the distance between successive anchors (i.e., peak-to-peak interval (PPI)) as $PPI^{k}=T_{1}^{k+1}-T_{1}^{k}$, as shown in Figure 1.
@@ II-A3 Problem Statement for Domain Transformation
In addition, the constant and abrupt noises heavily affect the quality of contactless cardiac monitoring and have been investigated a lot, especially for the algorithms in coarse cardiac feature monitoring [6, 17, 19]. However, the fine-grained ECG recovery could only realized by deep-learning methods, and the noise robustness of the deep-learning model has never been evaluated in the literature [2, 4, 3, 5]. Therefore, radarODE-MTL dissects the long-term ECG recovery into three tasks, and hence each decoder only focuses on extracting the cardiac feature with different granularity, aiming to improve the accuracy and noise robustness of the radar-based ECG recovery.
@@ II-B2 Problem Statement for Designing MTL Optimization Strategies
The existing methods perform not well on the proposed radarODE-MTL framework because most methods aim to treat all the tasks equally and pay too much attention to the easy tasks with the least achievement after convergence (e.g., slow learning rate in GradNorm [29], small singular value in Aligned-MTL [21]), while the hard task tolerates a slow convergence rate due to the limited gradient magnitudes or update frequencies [16]. Several studies in the literature proposed to increase the weight for the hard task metered the learning rate [16, 30]. However, the forcible change of the weight may aggravate the gradient conflict and hence degrade other tasks, because the loss-balancing method can not alleviate the gradient conflict issue [24].
@@ III-A Overview of radarODE-MTL with EGA Strategy
The aforementioned three deconstructed tasks for radar-based ECG recovery can be realized by the proposed radarODE-MTL framework as shown in Figure 3, and the dataset used for training and validation is provided in [2]. Firstly, the $50$ synchronous radar signals will be pre-processed into spectrograms by synchrosqueezed wavelet transform (SST) to highlight the central frequencies for locating the prominent vibrations $v_{1}$ and $v_{2}$. Then, radarODE-MTL is designed to generate the long-term ECG recovery in an end-to-end manner with certain shared layers to capture the common representations for all tasks and three task-specific decoders to recover the ECG morphological features, detect ECG anchors (R peaks) and estimate single-cardiac-cycle length respectively, as shown in Figure 3(a)-(d).
Figure 3: Overview of the radarODE-MTL framework with EGA strategy: (a) Shared backbone extracts time-frequency features from the signal spectrograms; (b) Morphological decoder only reconstructs the shape of the current ECG piece; (c) ECG anchor decoder estimates the time-index of anchors (R peaks); (d) Cycle length decoder estimates the length of the current cardiac cycle; (e) The proposed EGA strategy for optimizing shared parameter space.
@@ III-B Backbone and Encoder
The backbone of radarODE-MTL is used to extract the latent features from the input SST spectrograms and is expected to figure out the remarkable patterns for vibrations $v_{1}$ and $v_{2}$ with certain central frequencies and periodicity. Specifically, ResNet is adopted in this work as the backbone and has been proven to be an efficient structure in computer vision or signal processing [32, 33, 34]. Then, the encoder contains only one 2D convolutional layer to further compress the feature in the time-frequency domain into the 1D time domain for later processing. The performance has been verified in our previous work with the detailed structure shown in [5].
@@ III-F3 Eccentric Gradient Alignment
Methods Tasks ECG Shape Recovery Cycle Length Estimation ECG Anchor Estimation $\Delta m\%\uparrow$
RMSE (mV) $\downarrow$ PCC $\uparrow$ PPI Error (ms) $\downarrow$ Timing Error (ms) $\downarrow$ MDR $\downarrow$
Single-task baseline 0.106 86.6% 9.6 7.5 6.67% 0%
Loss Balancing Methods
Equal Weight 0.125 79.7% 8.0 9.7 5.51% -0.73%
UW [28] 0.066 88.5% 11.2 5.5 6.44% 5.95%
GLS [27] 0.087 87.3% 14.1 6.7 4.32% -4.65%
DWA [26] 0.133 80.7% 8.3 6.4 5.33% 4.84%
STCH [25] 0.070 88.0% 13.9 5.5 3.28% 3.90%
Gradient Balancing Methods
CAGrad [24] 0.107 84.2% 10.2 6.2 3.98% 6.72%
IMTL [23] 0.088 89.4% 9.3 6.0 6.22% 8.90%
MoCo [22] 0.179 61.0% 8.7 6.8 4.27% -5.72%
Aligned-MTL [21] 0.092 87.9% 10.0 6.9 3.52% 10.26%
EGA ($T=0.1$) 0.119 79.0% 10.6 6.8 3.34% 2.74%
EGA ($T=0.5$) 0.082 89.6% 9.9 6.3 4.19% 12.10%
EGA ($T=1.0$) 0.085 87.4% 8.5 7.2 4.31% 13.95%
EGA ($T=1.5$) 0.105 82.9% 8.1 6.3 5.13% 10.95%
EGA ($T=2.0$) 0.091 86.3% 9.2 7.3 4.01% 10.78%
@@ IV-A Dataset and Implementation
IV-A1 Dataset
MMECG [2] is a dataset used for radar-based ECG recovery and is collected by TI AWR-1843 radar with $77$Ghz start frequency and $3.8$GHz bandwidth. In addition, the respiration noise has been filtered in pre-processing, but this dataset still contains signals with low SNR or RBM noise.
@@ IV-A1 Dataset
IV-A2 Implementation Details
The proposed radarODE-MTL along with the radarODE [5] and MMECG [2] are coded using PyTorch and trained on the NVIDIA RTX A4000 ($16$GB) for $200$ epochs with batch size $32$, SGD optimizer [37], learning rate $5\times 10^{-3}$, weight decay $5\times 10^{-4}$ and momentum $0.937$. The dataset is split into training, validation and testing sets following $80\%:10\%:10\%$. At last, the Python package NeuriKit2 [9] is applied to all the evaluations regarding ECG signals, such as the identification of single cardiac cycles, PQRST peaks detection and heart rate estimation.
The deep learning framework used for scene understanding is implemented in [10] with many popular MTL optimization strategies embedded for comparison. The training is on the same GPU as before with $200$ epochs, batch size $18$, Adam optimizer [38], learning rate $10^{-4}$ and weight decay $10^{-5}$.
@@ IV-B1 Radar-based ECG Recovery
The performance of EGA is evaluated on three tasks in terms of different metrics: RMSE/PCC for the recovered single-cycle ECG pieces, absolute PPI Error for the cycle lengths estimation and absolute Timing Error and missed detected rate (MDR) for the anchors prediction, with the corresponding comparison across other MTL optimization strategies as shown in Table I. In addition, the last column $\Delta m\%$ in Table I shows a comprehensively assessment across $n$ tasks and is calculated as:
@@
where $n_{i}$ is the number of metrics for task $i$, $M_{m,i,j}$ means the performance of a method $m$ on the task $i$ measured with the metric $j$, $M_{b,i,j}$ represents the performance for the single-task baseline, and $S_{i,j}=1/0$ if lower/higher values are better for the current metric (indicated by $\downarrow$/$\uparrow$ / ↑).
In general, the proposed EGA strategy meets the expectation by adjusting the value of $T$ with the following evaluations:
  • EGA with $T=1.0$ achieves the largest improvement of $\Delta m\%$ but none of the individual metrics gets the best or second-best result, and $T=1.0$ can be viewed as a suitable estimation of intrinsic task difficulty, earning unbiased improvements on all tasks.
  • EGA with $T=0.5$ obtains the second-best overall improvement and becomes the best in learning ECG morphological features according to RMSE/PCC, indicating $T=0.5$ slightly overrates the difficulty of Task $1$.
@@
  • EGA with large $T$ values ($1.5$ and $2.0$) tend to evenly distribute the task difficulty weights, and the performance should be similar to other orthogonality-based method (e.g., Aligned-MTL).
@@
Method Segmentation $\uparrow$ Depth Estimation $\downarrow$ Surface Normal Prediction $\Delta m\%\uparrow$
@@
Single-task baseline 52.08 74.11 0.4147 0.1751 23.83 17.36 34.34 60.22 71.47 0.0%
@@
Equal Weight 53.36 74.94 0.3953 0.1672 24.35 17.55 34.22 59.64 70.71 1.74%
UW [28] 53.33 75.43 0.3878 0.1639 24.03 17.24 34.80 60.33 71.31 2.92%
GLS [27] 53.04 74.68 0.3951 0.1600 24.03 17.30 34.78 60.17 71.28 2.69%
DWA [26] 53.12 75.23 0.3883 0.1615 24.26 17.60 34.25 59.51 70.62 2.55%
STCH [25] 52.87 74.78 0.3915 0.1615 23.27 16.34 36.61 62.33 72.98 4.00%
@@
CAGrad [24] 52.19 74.07 0.3976 0.1634 23.83 17.16 34.89 60.65 71.77 2.09%
IMTL [23] 52.34 74.35 0.3897 0.1579 23.76 17.00 35.28 60.92 71.89 3.24%
MoCo [22] 52.78 74.59 0.3858 0.1612 23.34 16.51 36.21 61.90 72.65 3.94%
Aligned-MTL [21] 52.19 74.17 0.3911 0.1605 23.44 16.73 35.45 61.74 72.70 3.23%
EGA ($T=0.1$) 52.16 74.23 0.3944 0.1651 23.32 16.62 35.87 61.81 72.72 2.84%
EGA ($T=0.5$) 51.82 73.98 0.3904 0.1614 23.41 16.66 35.87 61.65 72.51 3.11%
EGA ($T=1.0$) 51.75 74.38 0.3913 0.1609 23.09 16.29 36.54 62.51 73.22 3.71%
EGA ($T=1.5$) 52.37 74.65 0.3950 0.1571 23.15 16.46 36.07 62.22 73.07 3.96%
EGA ($T=2.0$) 52.18 74.23 0.3922 0.1605 23.28 16.61 35.77 61.95 72.86 3.40%

### v2

@@ I Introduction
  • To the best of our knowledge, this is the first work that investigates the noise robustness in radar-based ECG recovery against constant or abrupt noise by modeling the cardiac domain transformation as three tasks. An end-to-end MTL framework named radarODE-MTL is accordingly proposed to realize these tasks and leverage adjacent cardiac cycles to compensate for the distorted one.
  • Sufficient experiments show that the proposed radarODE-MTL with EGA optimization strategy outperforms other frameworks and optimization strategies under various noise conditions and datasets, and the deconstructed tasks in radarODE-MTL could further improve the interpretability in radar-based ECG recovery.
@@ II-A1 Signal Model for Cardiac Mechanical Activities
In radar-based ECG recovery, the baseband signal is normally pre-processed using bandpass filter, differentiator and digital beamforming to remove the background and respiration noise to enhance cardiac-related features [6, 2, 17]. According to our previous work [5], the fine-grained cardiac mechanical activities include aortic valve opening/closure (AO/AC) and mitral valve opening/closure (MO/MC), revealed by the corresponding prominent vibrations $v_{1}$ and $v_{2}$ as measured in radar signal $x(t)$ as depicted in Figure 1. Therefore, the resultant radar signal $x(t)$ can be expressed for $K$ cardiac cycles as:
@@
where $a_{1}^{k}$, $b_{1}^{k}$ and $a_{2}^{k}$, $b_{2}^{k}$ jointly determine the amplitudes and lengths of the first and second prominent vibrations for $k^{th}$ cardiac cycle, $f_{1}^{k}$, $f_{2}^{k}$ are the corresponding central frequencies and $T_{1}^{k}$, $T_{2}^{k}$ represent when the vibrations happen. In addition, $n_{abr}(t)$ represents the abrupt noises (e.g., RBM) and $n_{con}(t)$ describes many other constant noises that affect the signal-to-noise ratio (SNR), such as thermal noise [6, 19], monitoring from random directions [20] and long-range monitoring [17].
@@ II-A2 Model of Domain Transformation
  • Task $1$: The reconstruction of the morphological features aims to map the single-cycle cardiac activities $x(t)$ to ECG with the deep neural network acting as a mapping function as $x_{ecg}(t)=\mathcal{T}(x(t))$.
@@
  • Task $3$: The prediction of the cardiac cycle length is equivalent to finding the peak-to-peak interval (PPI) used for resizing $x_{ecg}$ obtained in Task $1$, as shown in Figure 1.
Theoretically, PPI can be directly obtained from $\mathbf{R}$ as $T_{1}^{k+1}-T_{1}^{k}$, but it is necessary to reckon the PPI estimation to be an individual task in practice, because if one R peak fails to be detected in $\mathbf{R}$, the resultant PPI will be extremely large, destroying the long-term ECG recovery.
@@ II-A3 Problem Statement for Domain Transformation
In addition, the fine-grained ECG recovery could only realized by deep-learning methods, and the noise robustness of the deep-learning model has never been evaluated in the literature [2, 4, 3, 5]. Therefore, radarODE-MTL dissects the long-term ECG recovery into three tasks, and hence each decoder only focuses on extracting the cardiac feature with different granularity, aiming to improve the accuracy and noise robustness of the radar-based ECG recovery.
@@ II-B2 Problem Statement for Designing MTL Optimization Strategies
The existing methods perform not well on the proposed radarODE-MTL framework because most methods aim to treat all the tasks equally and pay too much attention to the easy tasks with the least achievement after convergence (e.g., slow learning rate in GradNorm [30], small singular value in Aligned-MTL [22]), while the hard task tolerates a slow convergence rate due to the limited gradient magnitudes or update frequencies [16]. Several studies in the literature proposed to increase the weight for the hard task metered the learning rate [16]. However, the forcible change of the weight may aggravate the gradient conflict and hence degrade other tasks, because the loss-balancing method can not alleviate the gradient conflict issue [25].
@@ III-A Overview of radarODE-MTL with EGA Strategy
The aforementioned three deconstructed tasks for radar-based ECG recovery can be realized by the proposed radarODE-MTL framework as shown in Figure 3, and the dataset used for training and validation is provided in [2]. Firstly, the $50$ synchronous radar signals will be pre-processed into spectrograms by synchrosqueezed transform (SST) to highlight the central frequencies for locating the prominent vibrations $v_{1}$ and $v_{2}$. Then, radarODE-MTL is designed to generate the long-term ECG recovery in an end-to-end manner with certain shared layers to capture the common representations for all tasks and three task-specific decoders to recover the ECG morphological features, detect ECG anchors (R peaks) and estimate single-cardiac-cycle length respectively, as shown in Figure 3(a)-(d).
Figure 3: Overview of the radarODE-MTL framework with EGA strategy: (a) Shared backbone extracts time-frequency features from the signal spectrograms with four layers of residual block; (b) Morphological decoder only reconstructs the shape of the current ECG piece; (c) ECG anchor decoder estimates the time-index of anchors (R peaks); (d) Cycle length decoder estimates the length of the current cardiac cycle; (e) The proposed EGA strategy for optimizing shared parameter space.
@@ III-B Backbone and Encoder
The backbone of radarODE-MTL is used to extract the latent features from the input SST spectrograms as shown in Figure 3(a) and is expected to figure out the remarkable patterns for vibrations $v_{1}$ and $v_{2}$ with certain central frequencies and periodicity. Specifically, four residual blocks are adopted in this work as the backbone because the ResNet has been proven to be an efficient structure in computer vision or signal processing [32, 33, 34]. Then, the encoder contains only one 2D convolutional layer to further compress the feature in the time-frequency domain into the 1D time domain for later processing. The performance of the backbone and encoder has been verified in our previous work with the detailed structure shown in [5].
@@ IV-A Dataset and Implementation
IV-A1 Dataset for ECG Recovery
MMECG [2] is a dataset used for radar-based ECG recovery and is collected by TI AWR-1843 radar with $77$GHz start frequency and $3.8$GHz bandwidth with the scenario of data collection shown in Figure 5. A total of $91$ trials for $11$ subjects ($8$ males and $3$ females) are included in the dataset, and each trial lasts for $3$ minutes with synchronous ECG/radar signals sampled at $200$Hz. Following the link budget analysis in [18, 37], the collected raw radar signals have a good SNR level of 37dB and are enough for later signal processing or model training.
All the subjects are healthy people without knowing diseases (e.g., premature ventricular contractions [4]) that may change the common ECG patterns and are asked to keep a quasi-static status to ensure good SNR with the least RBM noise. In addition, the ground ECG signal is collected by TI ADS1292 board with AC coupling and integrated right-leg drive (RLD) amplifier to remove potential baseline drift or power-line noise [38].
Figure 5: Scenario for data collection from quasi-static subject [2].
IV-A2 Dataset for Evaluating EGA
@@
IV-A3 Implementation Details
The proposed radarODE-MTL along with the radarODE [5] and MMECG [2] are coded using PyTorch and trained on the NVIDIA RTX A4000 ($16$GB) for $200$ epochs with SGD optimizer [39]. The hyperparameters used for training are empirically obtained as batch size $32$, learning rate $5\times 10^{-3}$, weight decay $5\times 10^{-4}$ and momentum $0.937$. The dataset is split based on different subjects, with the trials from $1$ fixed subject for testing and the other $10$ subjects alternatively selected for training or validation (i.e., $11$-fold cross-validation), ensuring to make use of all the possible trials while not involving the testing data in the training phase. At last, the Python package NeuriKit2 [9] is applied to all the evaluations regarding ECG signals, such as the identification of single cardiac cycles, PQRST peaks detection and heart rate estimation.
The deep learning framework used for scene understanding is implemented in [10] with many popular MTL optimization strategies embedded for comparison and optimal hyperparameters provided for training. The training is on the same GPU as before with $200$ epochs, batch size $18$, Adam optimizer [40], learning rate $10^{-4}$ and weight decay $10^{-5}$.
@@ IV-B1 Radar-based ECG Recovery
The performance of EGA is evaluated on three tasks in terms of different metrics: RMSE, PCC and coefficient of determination $R^{2}$ for the recovered single-cycle ECG pieces; absolute PPI Error for the cycle lengths estimation; and absolute Timing Error and missed detected rate (MDR) for the anchors prediction, with the corresponding comparison across other MTL optimization strategies as shown in Table I. In addition, all the experiments are repeated five times, and the last column $\Delta m\%$ in Table I shows a comprehensive assessment across $n$ tasks with $95\%$ confidence interval (CI) and is calculated as:
@@
where $n_{i}$ is the number of metrics for task $i$, $M_{m,i,j}$ means the performance of a method $m$ on the task $i$ measured with the metric $j$, $M_{b,i,j}$ represents the performance for the single-task baseline, and $S_{i,j}=1/0$ if lower/higher values are better for the current metric (indicated by $\downarrow$/$\uparrow$ / ↑). Lastly, the T-test is adopted with the $P$-value calculated for all the experiments as shown in Table I, and the statistical analysis will be given at the end of each subsection.
Methods Tasks ECG Shape Recovery Cycle Length Estimation ECG Anchor Estimation $\Delta m\%\uparrow$ $P$ Value (×$\times 10^{-2}$)
RMSE (mV) $\downarrow$ PCC $\uparrow$ $R^{2}$uparrow PPI Error (ms) $\downarrow$ Timing Error (ms) $\downarrow$ MDR $\downarrow$
Single-task baseline 0.106 86.6% 0.81 9.6 7.5 6.67% 0.00$\pm$1.43 -
Loss Balancing Methods
Equal Weight 0.125 79.7% 0.63 8.0 9.7 5.51% -1.78$\pm$2.16 9.26
UW [29] 0.066 88.5% 0.85 11.2 5.5 6.44% 4.04$\pm$3.79 2.43
GLS [28] 0.087 87.3% 0.81 14.1 6.7 4.32% -5.89$\pm$2.02 0.02
DWA [27] 0.133 80.7% 0.79 8.3 6.4 5.33% 6.45$\pm$3.71 0.20
STCH [26] 0.070 88.0% 0.86 13.9 5.5 3.28% 2.90$\pm$3.21 5.12
Gradient Balancing Methods
CAGrad [25] 0.107 84.2% 0.79 10.2 6.2 3.98% 6.84$\pm$2.12 0.01
IMTL [24] 0.088 89.4% 0.86 9.3 6.0 6.22% 8.43$\pm$1.39 0.00
MoCo [23] 0.179 61.0% 0.66 8.7 6.8 4.27% -2.32$\pm$1.37 1.16
Aligned-MTL [22] 0.092 87.9% 0.84 10.0 6.9 3.52% 10.14$\pm$1.11 0.00
EGA ($T=0.1$) 0.119 79.0% 0.72 10.6 6.8 3.34% 2.83$\pm$0.98 0.19
EGA ($T=0.5$) 0.082 89.6% 0.87 9.9 6.3 4.19% 11.55$\pm$1.44 0.00
EGA ($T=1.0$) 0.085 87.4% 0.85 8.5 7.2 4.31% 13.37$\pm$1.36 0.00
EGA ($T=1.5$) 0.105 82.9% 0.78 8.1 6.3 5.13% 10.94$\pm$1.30 0.00
EGA ($T=2.0$) 0.091 86.3% 0.84 9.2 7.3 4.01% 10.43$\pm$0.95 0.00
In general, the effect of unbalanced task-specific gradients is revealed by using equal weight as shown in Table I, and the performance of the hard task (ECG shape recovery) is much worse than baseline while the PPI error even achieves the best accuracy. After balancing the magnitudes and directions of task-specific gradients, the proposed EGA strategy meets the expectation by adjusting the value of $T$ with the following evaluations:
  • EGA with $T=1.0$ achieves the largest improvement with $\Delta m\%=13.37$ but none of the individual metrics gets the best or second-best result, and $T=1.0$ can be viewed as a suitable estimation of intrinsic task difficulty to achieve unbiased improvements across all tasks.
  • EGA with $T=0.5$ obtains the second-best overall performance with $\Delta m\%=11.55$ and becomes the best in learning ECG morphological features according to RMSE/PCC/$R^{2}$, indicating $T=0.5$ slightly overrates the difficulty of Task $1$.
@@
  • EGA with large $T$ values ($1.5$ and $2.0$) tend to evenly distribute the task difficulty weights, and the performance should be similar to other orthogonality-based methods (e.g., Aligned-MTL).
@@
Method Segmentation $\uparrow$ Depth Estimation $\downarrow$ Surface Normal Prediction $\Delta m\%\uparrow$ $P$ Value (×$\times 10^{-2}$)
@@
Single-task baseline 52.08 74.11 0.4147 0.1751 23.83 17.36 34.34 60.22 71.47 0.00$\pm$0.19 -
@@
Equal Weight 53.36 74.94 0.3953 0.1672 24.35 17.55 34.22 59.64 70.71 1.75$\pm$1.60 1.67
UW [29] 53.33 75.43 0.3878 0.1639 24.03 17.24 34.80 60.33 71.31 2.92$\pm$2.05 0.44
GLS [28] 53.04 74.68 0.3951 0.1600 24.03 17.30 34.78 60.17 71.28 2.69$\pm$1.52 0.12
DWA [27] 53.12 75.23 0.3883 0.1615 24.26 17.60 34.25 59.51 70.62 2.55$\pm$1.91 0.62
STCH [26] 52.87 74.78 0.3915 0.1615 23.27 16.34 36.61 62.33 72.98 3.99$\pm$0.61 0.00
@@
CAGrad [25] 52.19 74.07 0.3976 0.1634 23.83 17.16 34.89 60.65 71.77 2.09$\pm$1.11 0.09
IMTL [24] 52.34 74.35 0.3897 0.1579 23.76 17.00 35.28 60.92 71.89 3.24$\pm$0.78 0.00
MoCo [23] 52.78 74.59 0.3858 0.1612 23.34 16.51 36.21 61.90 72.65 3.94$\pm$0.72 0.00
Aligned-MTL [22] 52.19 74.17 0.3911 0.1605 23.44 16.73 35.45 61.74 72.70 3.24$\pm$1.08 0.00
EGA ($T=0.1$) 52.16 74.23 0.3944 0.1651 23.32 16.62 35.87 61.81 72.72 2.84$\pm$0.88 0.00
EGA ($T=0.5$) 51.82 73.98 0.3904 0.1614 23.41 16.66 35.87 61.65 72.51 3.11$\pm$0.51 0.00
EGA ($T=1.0$) 51.75 74.38 0.3913 0.1609 23.09 16.29 36.54 62.51 73.22 3.71$\pm$1.08 0.00
EGA ($T=1.5$) 52.37 74.65 0.3950 0.1571 23.15 16.46 36.07 62.22 73.07 3.96$\pm$0.98 0.00
EGA ($T=2.0$) 52.18 74.23 0.3922 0.1605 23.28 16.61 35.77 61.95 72.86 3.39$\pm$1.18 0.00
@@
Figure 6: Illustration of the recovered ECG with ground truth (GT) under different frameworks and noises: (a) - (d) Radar signal with good quality and the recovered ECG signals; (e) - (h) Radar signal with constant noise and the recovered ECG signals; (i) - (l) Radar signal with abrupt noise and the recovered ECG signals.